On January 1, 2016, Halstead, Inc., purchased 84,000 shares of Sedgwick Company common stock for $1,499,000, giving Halstead 25 percent ownership and the ability to apply significant influence over Sedgwick. Any excess of cost over book value acquired was attributed solely to goodwill. Sedgwick reports net income and dividends as follows. These amounts are assumed to have occurred evenly throughout these years. Dividends are declared and paid in the same period. Annual Cash Dividends (pai quarterly) $122,ee8 128,000 167,000 Net Income 2016 2017 2018 $366,00e 411,000 617,8ee On July 1, 2018, Halstead sells 12,800 shares of this investment for $29 per share, thus reducing its interest from 25 to 21 percent, but maintaining its significant influence. Determine the amounts that would appear on Halsteads 2018 income statement relating to its ownership and partial sale of its investment in Sedgwicks common stock. As total income accrual (no unearned gains) As on sale of shares
Answer #1

Solution:
Halstead Inc.

Particulars Amount ($)
Equity Method Income Accrual (2018)
(25%*617,000*1/2 year) $77,125
(21%*617,000*1/2 year) $64,785
Total Income Accrual $141,910
Gain on sale (Working Note) $91,218
Total Effect on income $233,128

Working Note:
Cost of acquisition = $1,499,000
2016 Income accrual (366,000*25%) = $91,500
2016 dividend (122,000*25%) = ($30,500)
2017 Income accrual (411,000*25%) = $102,750
2017 dividend (128,000*25%) = ($32,000)
2018 Income accrual (617,000*25%*1/2) = $77,125
2018 dividend (167,000*25%*1/2) = ($20,875)
Book value of 77,125 shares as on July 1, 2018 = $1,687,000
Cash proceeds from sale (12,800*29) = $371,200
Less: Book value of shares sold (1,687,000/77,125)*12,800 = $279,982
Gain on Sale = $91,218